OpenStudy (anonymous):

Solving system by graphing y=x+6 y=-3x+6

OpenStudy (anonymous):

How do you fo this

OpenStudy (anonymous):

what are you trying to find?

OpenStudy (anonymous):

Are you trying to figure where the lines intercept at or the x value?

OpenStudy (anonymous):

We have to have 2 line on the graph

OpenStudy (anonymous):

I'm assuming when you say solving system, you mean systems of equations by graphing.

OpenStudy (anonymous):

type y= for both equations and then graph or 2nd graph if its helpful

OpenStudy (anonymous):

You can easily just use substitution, so x+6 = -3x+6. solve and that will be your X.

OpenStudy (anonymous):

Yes systems of equations by graphing. I dont understand

OpenStudy (anonymous):

What dont you understand ? graphing?

OpenStudy (anonymous):

How to solve systems y graphing

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

Do you have GeoAlgebra on your computer?

OpenStudy (anonymous):

Will help with the graphing.

OpenStudy (anonymous):

Do you have a graphing calculator? or online?

OpenStudy (triciaal):

I can help you without the calculator or online. to solve by graphing means just that. plot points get the lines. the solution is where the lines intersect. to get the line you need at least 2 points. the easiest to use when x = 0 what is y and when y = 0 what is x

OpenStudy (triciaal):

let y=x+6 represent line 1 let y=-3x+6 represent line 2

OpenStudy (triciaal):

line 1 when x =0 y = 6 when y = 0 x = -6 you have points (0, 6) and (-6, 0) line 2 when x = 0 y = 6 and when y = 0 x = 2 points are (0, 6) and (2, 0)

OpenStudy (triciaal):